OpenStudy (anonymous):

I will give medals Simplify:

OpenStudy (anonymous):

\[\sqrt{36/121}\]

OpenStudy (anonymous):

6/11

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

one sec, lemme check :)

Nnesha (nnesha):

\[\huge\rm \sqrt{\frac{ a }{ b }} = \frac{ \sqrt{a} }{ \sqrt{b} }\]

OpenStudy (anonymous):

it looks like WorkHelp is correct

OpenStudy (anonymous):

yes he is.

OpenStudy (anonymous):

kk, thanks ^.^

Nnesha (nnesha):

so \[\sqrt{\frac{ 36 }{ 121 }} \] can be written as \[ \frac{ \sqrt{36} }{ \sqrt{121} }\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

glad i could help